# Fair Concurrent Training of Multiple Models in Federated Learning

### v1

@@ V-B Fairness with Respect to Max-Min Task Take-Up Rates
Suppose that all user bids for a given task $s$ are independent samples from a probability distribution $f_{s}$. Let $x_{s}\sim\bar{f}_{s}$ denote the minimum of the $K$ user bids for task $s$. Then the max-min fair auction has a smaller probability that at least one task is allocated no users compared to the budget-fair auction: $\mathbb{P}\left[\sum_{s=1}^{S}x_{s}\geq B\right]\leq\mathbb{P}\left[\max_{s=1}^{S}x_{s}\geq\frac{B}{S}\right].$ For example, if $f_{s}\sim\exp(\lambda)$ and $S=2$, then the probability at least one task is allocated no users in the max-min fair auction is $\exp(-B\lambda)(1+B\lambda)$, which is strictly greater than the probability at least one task is allocated no users in the budget-fair auction, $\exp(-B\lambda)\left(2\exp\left(\frac{B\lambda}{2}\right)-1\right)$ if $\lambda B>0$.
@@ V-C Convergence Guarantees of our Incentive-Aware Auctions
Under the Budget-Fair Auction, the probability $p_{B,b_{i,s}}^{join}$ that a user with bid $b_{i}$ joins task $s$, given a total budget of $B$ is $\sum_{\overline{t}}P(b_{i,s}\ \text{is}\ \overline{t}\text{-th smallest})P(\frac{B}{s\overline{t}}>b_{i,s})$. Under the Max-min Fair Auction, the probability $p^{join}_{B,b_{i,s}}$ is $\sum_{\overline{t}}P(b_{i,s}\text{is}\ \overline{t}\text{-th smallest)}P(\frac{B}{s\overline{t}}>b_{i,s})+\sum_{\hat{t}}P(b_{i,s}\text{is}\ \hat{t}\text{-thsmallest)}P(\frac{B}{s\hat{t}}<b_{i,s})P(\sum_{s\in S_{\hat{t}-1}}\underline{b}_{\hat{t},s}<C)$ for full participation, and $\sum_{\overline{t}}P(b_{i,s}\ \text{is}\ \overline{t}\text{-th smallest})P(\frac{B}{s\overline{t}}<b_{i,s})\times P(\sum_{s\in S_{\overline{t}-1}}\underline{b}_{\overline{t},s}>C)\min(1,\frac{C}{|S_{t-1}|b_{i,s}})$ for partial participation. Letting $\mathbb{W}$ be the set of auction winners, and $W_{K}$ be all the subsets of $K$ clients recruited from $N$ users:
@@ VI Evaluation
Figure 2: Results of Experiment 1: 3 tasks with varying difficulty levels. The 3 tasks have datasets MNIST, CIFAR-10, and Fashion-MNIST. FedFairMMFL achieves a higher minimum accuracy across tasks, as compared to the baselines, while maintaining the same or higher accuracy for the other 2 tasks (see Technical Report [36]).
@@
(c) The minimum accuracy after combining our auction strategies and baselines, with FedFairMMFL.
@@
Figure 5: Results of Experiment 4: Client Take-Up Rates. (NB= No Budget, NT= Non Truthful Mechanism). Our near-truthful auction MMFL Max-Min Fair (a) leads to a lower difference in the number of users across tasks, and (b) incentivizes more clients than any incentivization algorithm but the un-truthful GMMFair. Results of Experiments 5: How auctions impact learning. Fig. (c) shows the minimum accuracy after combining the results of our auction strategies and baselines, with FedFairMMFL. The MMFL Max-Min Fair Auction leads to higher accuracies.
@@
Finally we evaluate FedFairMMFL and our auction schemes. Our main results are that (i) FedFairMMFL consistently performs fairly across different experiment settings and (ii) our max-min auction increases the minimum learning accuracy across tasks in the budget constrained setting. We concurrently train (i) three tasks of varied difficulty levels, and evaluate the algorithms’ performance across varied numbers of (ii) tasks and (iii) clients. We analyze (iv) the impact of our auctions and baselines on the take-up rate over tasks, and analyze (v) the impact of these take-up rates on the tasks’ learning outcomes. We use $\alpha=3$ for FedFairMMFL in all experiments. The effect of different $\alpha$ is also measured (see Technical Report [36]).
@@
Our client-task allocation algorithm FedFairMMFL is measured against two baselines: a) Random, where clients are randomly allocated a task with equal probability each global epoch (note that this corresponds to taking $\alpha=1$ in our $\alpha$-fair allocation), and b) Round Robin [8], where active clients are allocated to tasks in a round-robin manner, sequentially across global epochs. Our main fairness metrics are the minimum test accuracy across all tasks and the variance of the test accuracies (cf. Lemma 1). As some models may have larger training losses by nature, e.g., if they are defined on different datasets, making them more easily favored by the algorithm, we use the test accuracy instead of training loss in FedFairMMFL. Each experiment’s results are averaged over 3-5 random seeds - we plot the average values (dotted line), along with the maximum and minimum values indicated by the shaded area. Each client’s data is sampled from either all classes (i.i.d.) or a randomly chosen half of the classes (non-i.i.d.).
@@
Experiment 1: Tasks of Varying Difficulty Levels. To model differing task difficulties, our tasks are to train image recognition models on MNIST and Fashion-MNIST on a convolutional neural network with 2 convolutional, 2 pooling, and 2 fully connected layers; and CIFAR-10 with a pre-activated ResNet. Experiments are run with 120 clients having $150-250$ datapoints from five (randomly chosen) of the ten classes (i.e., non i.i.d. data); $35\%$ of clients, chosen uniformly at random, are active each round. As seen in Fig. 2 a), FedFairMMFL has a higher minimum accuracy than the random and round robin baselines. FedFairMMFL adaptively allocates more clients to the prevailing worst-performing task, resulting in a higher accuracy for this task (Task 3: Fashion-MNIST), as seen in Fig. 2 b), while maintaining the same or a slightly higher accuracy for the other tasks (see Tech Report[36]). Though we might guess that CIFAR-10 would be the worst-performing task, as it uses the largest model, FedFairMMFL identifies that Fashion-MNIST is the worst-performing task, and it achieves fair convergence even in this challenging setting where Fashion-MNIST persistently performs worse than the other tasks.
@@
Experiment 2: Increasing the number of tasks. To further understand the various algorithms (FedFairMMFL and the random allocation and round robin baselines) under different settings, their performance is measured as we increase the number of tasks, from 3 to 4, 5, 6 and 10. The tasks include MNIST, Fashion-MNIST, CIFAR-10, and EMNIST. We use the same networks as in Experiment 1 and a similar convolutional neural network with 2 convolutional, 2 pooling, and 2 fully connected layers for the EMNIST task. Experiments are run with 20 clients, having a participation rate of 1. Each client has $400-600$ datapoints from each dataset for each task. We plot the variance amongst tasks on average, as well as at global epoch round numbers 10, 60, and 120. As seen in Fig. 3, with the increase in the round number, all algorithms’ variance decreases, and FedFairMMFL has the lowest variance for all settings. This can be explained through the fact that FedFairMMFL dynamically allocates clients with a higher probability to tasks which are currently performing worse, eliminating the difference among tasks as training progresses. This effect remains even as we increase the number of tasks, which slows the training for all tasks as the clients are being shared among more tasks - see Fig. 3 (e), (f), and (g), which compare the minimum accuracy among different algorithms when task numbers are 3, 6, and 10. FedFairMMFL converges faster than both baselines.
@@
Experiment 4: Client Take-Up Rates. We compare the number of clients incentivized to train each task under our budget-fair (Section V-A) and max-min fair (Section V-B) auctions, as well as the (un-truthful) GMMFair algorithm, which upper-bounds the max-min fair auction result, and 4 other baselines: valThreshold 0.4 and valThreshold 0.6 assume no budget but that all users with cost (valuation) $c_{i,s}<0.4$ (respectively $c_{i,s}<0.6$) will agree to train the task; these simulate posted-price mechanisms that show the effect of the incentivization budget. Greedy Within Budget and Random Within Budget evenly divide the budget across tasks. Clients are added by ascending bid (for Greedy Within Budget) and randomly (for Random Within Budget), until the budget is consumed; comparison to these baselines shows the cost of enforcing truthfulness in the budget-fair auction. We consider two tasks and 100 clients; client bids are between 0 and 1 and are independently drawn from a truncated Gaussian distribution for Task 1 and an increasing linear distribution for Task 2. Results are averaged across 5 random seeds.
@@
Experiment 5: Take-Up Rates and Learning Outcomes. We finally consider a constrained budget ($B=29$) and evaluate how the outcome of the client recruitment strategies impact MMFL model training. Fig. 5 c) shows the minimum test accuracy across the two tasks (MNIST and CIFAR-10) at each training round, using FedFairMMFL for client-task allocation. GMMFair consistently has the highest minimum accuracy across tasks, likely because it recruits the most clients to each task, as shown in Experiment 4 (Fig. 5 b). Our MMFL auction, which additionally has the benefit of near-truthfulness, achieves nearly the same minimum accuracy, showing that it performs well in tandem with FedFairMMFL.
@@ VII Conclusion and Future Directions
In this paper, we consider the fair training of concurrent MMFL tasks. Ours is the first work to both theoretically and experimentally validate algorithms that ensure fair outcomes across different learning tasks. We first propose FedFairMMFL, a simple algorithm to allocate clients to tasks in each training round, showing that it permits theoretical convergence guarantees. We then propose two auction mechanisms, Budget-Fair and MMFL Max-Min Fair, that seek to incentivize a fair number of clients to commit to training each task. Our experimental results for training different combinations of FL tasks show that FedFairMMFL and MMFL Max-Min Fair can jointly ensure fair convergence rates and final test accuracies across the different tasks. As the first work to consider fairness in the MMFL setting, we expect our work to seed many future directions, including how to balance achieving fairness over tasks with fairness over clients’ utilities. Another direction is to investigate fairness when tasks have differing difficulties and clients have heterogeneous, stochastic available computing resources for training each task.

### v2

@@ V-B Fairness with Respect to Max-Min Task Take-Up Rates
Suppose that all user bids for a given task $s$ are independent samples from a probability distribution $f_{s}$. Let $x_{s}\sim\bar{f}_{s}$ denote the minimum of the $K$ user bids for task $s$. Then the max-min fair auction has a smaller probability that at least one task is allocated no users compared to the budget-fair auction:
$\mathbb{P}\left[\sum_{s=1}^{S}x_{s}\geq B\right]\leq\mathbb{P}\left[\max_{s=1}^{S}x_{s}\geq\frac{B}{S}\right].$
For example, if $f_{s}\sim\exp(\lambda)$ and $S=2$, then the probability at least one task is allocated no users in the max-min fair auction is $\exp(-B\lambda)(1+B\lambda)$, which is strictly greater than the probability at least one task is allocated no users in the budget-fair auction, $\exp(-B\lambda)\left(2\exp\left(\frac{B\lambda}{2}\right)-1\right)$ if $\lambda B>0$.
@@ V-C Convergence Guarantees of our Incentive-Aware Auctions
Under the Budget-Fair Auction, the probability $p_{B,b_{i,s}}^{join}$ that a user with bid $b_{i}$ joins task $s$, given a total budget of $B$ is
$\sum_{\overline{t}}P(b_{i,s}\ \text{is}\ \overline{t}\text{-th smallest})P(\frac{B}{s\overline{t}}>b_{i,s}).$
Under the Max-min Fair Auction, the probability $p^{join}_{B,b_{i,s}}$ that a user with bid $b_{i}$ joins task $s$, given a total budget of $B$ is
$\sum_{\overline{t}}P(b_{i,s}\text{is}\ \overline{t}\text{-th smallest)}P(\frac{B}{s\overline{t}}>b_{i,s})$
$+$ $\sum_{\hat{t}}P(b_{i,s}\text{is}\ \hat{t}\text{-th smallest)}P(\frac{B}{s\hat{t}}<b_{i,s})P(\sum_{s\in S_{\hat{t}-1}}\underline{b}_{\hat{t},s}<C)$
for full participation, and
$\sum_{\overline{t}}P(b_{i,s}\ \text{is}\ \overline{t}\text{-th smallest})P(\frac{B}{s\overline{t}}<b_{i,s})$
$\times$ $P(\sum_{s\in S_{\overline{t}-1}}\underline{b}_{\overline{t},s}>C)\min(1,\frac{C}{|S_{t-1}|b_{i,s}})$
for partial participation. Letting $\mathbb{W}$ be the set of auction winners, and $W_{K}$ be all the subsets of $K$ clients recruited from $N$ users, we present a new convergence bound which takes into account the joint probability of users’ being in the auction’s winners set.
@@ VI Evaluation
(c) Accuracy for Task 1: MNIST.
(d) Accuracy for Task 2: CIFAR-10.
Figure 2: Results of Experiment 1: 3 tasks with varying difficulty levels. The 3 tasks have datasets MNIST, CIFAR-10, and Fashion-MNIST. FedFairMMFL achieves a higher minimum accuracy across tasks, as compared to the baselines, while maintaining the same or higher accuracy for the other 2 tasks.
@@
(a) Minimum accuracy over 3 tasks, MNIST, CIFAR-10, and Fashion-MNIST.
(b) Minimum accuracy over 4 tasks, MNIST, CIFAR-10, EMNIST, and Fashion-MNIST.
Figure 4: Comparison of two alpha-fair based methods. a) 3 tasks include MNIST, CIFAR-10, and Fashion-MNIST. b) 4 tasks include MNIST, CIFAR-10, EMNIST, and Fashion-MNIST. EMNIST convergence is extremely slow with q-Fel. FedFairMMFL converges faster and achieves a higher minimum accuracy across tasks compared to q-Fel for task fairness.
@@
(a) Random allocation.
(b) Round robin allocation.
(c) FedFairMMFL ($\alpha=1$).
(d) FedFairMMFL ($\alpha=2$).
(e) FedFairMMFL ($\alpha=3$).
(f) FedFairMMFL ($\alpha=4$).
(g) FedFairMMFL ($\alpha=5$).
(h) FedFairMMFL ($\alpha=6$).
Figure 6: Experiment 4 on different values of $\alpha$: (a)-(h) show the maps of clients allocation to 5 tasks for 120 rounds of training in two baseline algorithms and FedFairMMFL with different $\alpha$. (c) $\alpha=1$ yields a similar result as the (a) random allocation, and as $\alpha$ increases, more resources are allocated to the hardest task (task 3). The 5 tasks are MNIST, CIFAR-10, Fashion-MNIST, EMNIST, and CIFAR-10, respectively.
@@
(c) The minimum accuracy after combining our auction strategies and baselines, with FedFairMMFL, i.e. combining the client recruitment part and model training parts of the MMFL pipeline.
@@
Figure 7: Results of Experiment 5: Client Take-Up Rates. (NB= No Budget, NT= Non Truthful Mechanism). Our near-truthful auction MMFL Max-Min Fair (a) leads to a lower difference in the number of users across tasks, and (b) incentivizes more clients than any incentivization algorithm but the un-truthful GMMFair. Results of Experiments 6: How auctions impact learning. Fig. (c) shows the minimum accuracy after combining the results of our auction strategies and baselines, with FedFairMMFL. The MMFL Max-Min Fair Auction leads to higher accuracies.
@@
Finally we evaluate FedFairMMFL and our auction schemes. Our main results are that (i) FedFairMMFL consistently performs fairly across different experiment settings, increasing the minimum accuracy and decreasing the variance over tasks, while maintaining the average accuracy over tasks; and (ii) our max-min auction increases the minimum learning accuracy across tasks in the budget constrained setting. We concurrently train (i) three tasks of varied difficulty levels, and evaluate the algorithms’ performance across varied numbers of (ii) tasks and (iii) clients. We analyze (iv) the impact of our auctions and baselines on the take-up rate over tasks, and analyze (v) the impact of these take-up rates on the tasks’ learning outcomes. We use $\alpha=3$ for FedFairMMFL in all experiments. The effect of different $\alpha$ is also measured and discussed.
@@
Our client-task allocation algorithm FedFairMMFL is measured against three baselines: a) Random, where clients are randomly allocated a task with equal probability each global epoch (note that this corresponds to taking $\alpha=1$ in our $\alpha$-fair allocation), b) Round Robin [13], where active clients are allocated to tasks in a round-robin manner, sequentially across global epochs, and c) q-Fel [23], which also uses alpha-fairness to enhance fairness (but across clients and not tasks), where the parameter $q$ is the $\alpha$-equivalent. Instead of changing the sampling probability as in our FedFairMMFL, q-Fel directly changes the aggregation coefficients based on the $\alpha$-fairness function and adjusts the global learning rate accordingly. For the baseline, we use an adapted version of q-Fel, where fairness is over tasks and not clients. Our main fairness metrics are the minimum test accuracy across all tasks and the variance of the test accuracies (cf. Lemma 1). As some models may have larger training losses by nature, e.g., if they are defined on different datasets, making them more easily favored by the algorithm, we use the test accuracy instead of training loss in FedFairMMFL. Each experiment’s results are averaged over 3-5 random seeds - we plot the average values (dotted line), along with the maximum and minimum values indicated by the shaded area. Each client’s data is sampled from either all classes (i.i.d.) or a randomly chosen half of the classes (non-i.i.d.).
@@
Experiment 1: Tasks of Varying Difficulty Levels. To model differing task difficulties, our tasks are to train image recognition models on MNIST and Fashion-MNIST on a convolutional neural network with 2 convolutional, 2 pooling, and 2 fully connected layers; and CIFAR-10 with a pre-activated ResNet. Our algorithm does not assume that the task with larger model is the more difficult task, rather it takes the current performance (accuracy) levels as an indicator on which task needs more clients currently. Experiments are run with 120 clients having $150-250$ datapoints from five (randomly chosen) of the ten classes (i.e., non i.i.d. data); $35\%$ of clients, chosen uniformly at random, are active each round. As seen in Fig. 2 a), FedFairMMFL has a higher minimum accuracy than the random and round robin baselines. FedFairMMFL adaptively allocates more clients to the prevailing worst-performing task, resulting in a higher accuracy for this task (Task 3: Fashion-MNIST), as seen in Fig. 2 b), while maintaining the same or a slightly higher accuracy for the other tasks (see Fig. 2 c and d.) Though we might guess that CIFAR-10 would be the worst-performing task, as it uses the largest model, FedFairMMFL identifies that Fashion-MNIST is the worst-performing task, and it achieves fair convergence even in this challenging setting where Fashion-MNIST persistently performs worse than the other tasks.
TABLE II: Experiment 2 Quantitative Results. Our algorithm FedFairMMFL improves the minimum accuracy over tasks; at the same time, our average accuracy across tasks, is comparable with the baselines.
Methods 6 tasks 10 tasks
Avg Acc Min Acc Avg Acc Min Acc
Random 0.789 0.450 0.748 0.386
Round Robin 0.782 0.453 0.758 0.406
FedFairMMFL 0.783 0.475 0.754 0.452
@@
Experiment 2: Increasing the number of tasks. To further understand the various algorithms (FedFairMMFL and the random allocation and round robin baselines) under different settings, their performance is measured as we increase the number of tasks, from 3 to 4, 5, 6, and 10. The tasks include MNIST, Fashion-MNIST, CIFAR-10, and EMNIST. We use the same networks as in Experiment 1 and a similar convolutional neural network with 2 convolutional, 2 pooling, and 2 fully connected layers for the EMNIST task. Experiments are run with 20 clients, having a participation rate of 1. Each client has $400-600$ datapoints from each dataset for each task. We plot the variance amongst tasks on average, as well as the variance at global epoch round numbers 10, 60, and 120. As seen in Fig. 3, with the increase in the round number, all algorithms’ variance decreases, and FedFairMMFL has the lowest variance for all settings. This can be explained through the fact that FedFairMMFL dynamically allocates clients with a higher probability to tasks which are currently performing worse, eliminating the difference among tasks as training progresses. This effect remains even as we increase the number of tasks, which slows the training for all tasks as the clients are being shared among more tasks - see Fig. 3 (e), (f), and (g), which compare the minimum accuracy among different algorithms when task numbers are 3, 6, and 10. FedFairMMFL converges faster than both baselines. Table II shows the final accuracies for these algorithms. In terms of task fairness, FedFairMMFL achieves higher minimum accuracy across tasks. Besides, even though FedFairMMFL allocates more clients to tasks with the prevailing higher loss values during training, the average accuracy is not impacted much, and is comparable with the baselines, as seen in Table II.
Fig. 4 compares two alpha-fair based methods: our algorithm FedFairMMFL, and an adapted version of q-Fel [23], where fairness is over tasks and not clients. We notice that q-Fel converges quite slowly when we include EMNIST in the set of tasks. This is because, to guarantee convergence, q-Fel adjusts the aggregation coefficients and learning rate. For specific tasks, the adjusted global learning rate (which is inferred from the upper bound of the local Lipschitz constants of the gradients) can be extremely small, therefore slowing down convergence, which [23] also showed in their results. FedFairMMFL directly modifies the client sampling probability and does not modify the aggregation coefficients, providing a more reliable performance in diverse task settings.
@@
Experiment 4: The impact of different values of $\alpha$. To understand the effect of $\alpha$ on client-task allocation over the whole training process, we test the FedFairMMFL with multiple values of $\alpha$. Experimental results are shown in Fig. 6. Here all experiments have 5 tasks (MNIST, CIFAR-10, Fashion-MNIST, EMNIST, and CIFAR-10), and we have $20$ clients that participate in each round. Each client has $400-600$ datapoints from each dataset for each task. We use the same models as we used in Experiment 2 and 3. In this setting, EMNIST (task 3) is always the hardest task, and therefore gets more clients as $\alpha$ increases. When $\alpha=1$, allocation is uniform as the algorithm degenerates to the random client-task allocation. When $\alpha$ is too high, most of the resources will be allocated to only one task as shown in Fig. 6 (g)(h). This uneven allocation can lead to other simpler tasks not receiving sufficient training and requiring additional rounds to achieve satisfactory accuracy. Since $\alpha=3$ balances the fairness (in terms of the minimum accuracy and variance) and the training efficiency (in terms of the average accuracy) in a desirable way, we use $\alpha=3$ in all of our other experiments.
Experiment 5: Client Take-Up Rates. We compare the number of clients incentivized to train each task under our budget-fair (Section V-A) and max-min fair (Section V-B) auctions, as well as the (un-truthful) GMMFair algorithm, which upper-bounds the max-min fair auction result, and 4 other baselines: valThreshold 0.4 and valThreshold 0.6 assume no budget but that all users with cost (valuation) $c_{i,s}<0.4$ (respectively $c_{i,s}<0.6$) will agree to train the task; these simulate posted-price mechanisms that show the effect of the incentivization budget. Greedy Within Budget and Random Within Budget evenly divide the budget across tasks. Clients are added by ascending bid (for Greedy Within Budget) and randomly (for Random Within Budget), until the budget is consumed; comparison to these baselines shows the cost of enforcing truthfulness in the budget-fair auction. We consider two tasks and 100 clients; client bids are between 0 and 1 and are independently drawn from a truncated Gaussian distribution for Task 1 and an increasing linear distribution for Task 2. Results are averaged across 5 random seeds.
@@
Experiment 6: Take-Up Rates and Learning Outcomes. We finally consider a constrained budget ($B=29$) and evaluate how the outcome of the client recruitment strategies (our auction strategies and baselines) impact MMFL model training, i.e. combining the client recruitment and model training parts of the pipeline (Fig. 1). Specifically, the client recruitment strategies (MMFL Max-Min Fair, GMM Fair, and baselines valThreshold, Greedy Within Budget) determines which client has been recruited for which task. These clients will then be involved in MMFL training under our algorithm FedFairMMFL. Fig. 7 c) shows the minimum test accuracy across the two tasks (MNIST and CIFAR-10) at each training round, using FedFairMMFL for client-task allocation. GMMFair consistently has the highest minimum accuracy across tasks, likely because it recruits the most clients to each task, as shown in Experiment 5 (Fig. 7 b). Our MMFL auction, which additionally has the benefit of near-truthfulness, achieves nearly the same minimum accuracy, showing that it performs well in tandem with FedFairMMFL.
@@ VII Conclusion and Future Directions
In this paper, we consider the fair training of concurrent MMFL tasks. Ours is the first work to both theoretically and experimentally validate algorithms that ensure fair outcomes across different learning tasks, which are concurrently trained in an FL setting. We first propose FedFairMMFL, an algorithm to allocate clients to tasks in each training round, based on prevailing performance levels, showing that it permits theoretical convergence guarantees. We then propose two auction mechanisms, Budget-Fair and MMFL Max-Min Fair, that seek to incentivize a fair number of clients to commit to training each task. Our experimental results for training different combinations of FL tasks show that FedFairMMFL and MMFL Max-Min Fair can jointly ensure fair convergence rates and final test accuracies across the different tasks. As the first work to consider fairness in the MMFL setting, we expect our work to seed many future directions, including how to balance achieving fairness over tasks with fairness over clients’ utilities. Another direction is to investigate fairness when tasks have differing difficulties and clients have heterogeneous, and dynamically available computing resources for training.